Draw the structural formula of (3E,5Z)-5-ethyl-3,5-nonadiene.
Answer #1
Concepts and reason

The root word in the name of the compound indicates the number of carbon atoms in the longest carbon chain. The numbers before the name of the substituents and double or triple bonds represent the position of the substituents or double or triple bonds.

The E-Z system of nomenclature is used for alkenes having two different groups on the double bonded carbon atoms.

Fundamentals

CahnIngoldPrelogpriorityrulestowritethepriorityforthetwogroupsattachedtoadoublebondedcarbonatomareasfollows:1.Givethehighestprioritytotheatomwithhighestatomicnumberwhichisdirectlyattachedtochiralcarbon.2.Considerthedoublebondsortriplebondsastwoorthreesinglebonds.3.Iftwoatomswhicharedirectlyattachedtothechiralcarbonaresame,thenconsiderthenextsetofatomswhichisdirectlyattachedtothefirstatom.4.Iftwoisotopesareattachedtothechiralcarbon,thengivethehighestprioritytotheelementwithhighestmassnumber.\begin{array}{l}\\{\rm{Cahn - Ingold - Prelog priority rules to write the priority for the two groups attached }}\\\\{\rm{to a double}}\,\,{\rm{bonded carbon}}\,\,{\rm{atom}}\,\,{\rm{are}}\,\,{\rm{as}}\,\,{\rm{follows:}}\\\\{\rm{1}}{\rm{. Give the highest priority to the atom with highest atomic number which is directly }}\\\\{\rm{attached to chiral carbon}}{\rm{.}}\\\\{\rm{2}}{\rm{. Consider the double bonds or triple bonds as two or three single bonds}}{\rm{.}}\\\\{\rm{3}}{\rm{. If two atoms which are directly attached to the chiral carbon are same, then consider the }}\\\\{\rm{next set of atoms which is directly attached to the first atom}}{\rm{. }}\\\\{\rm{4}}{\rm{. If two isotopes are attached to the chiral carbon, then give the highest priority to the }}\\\\{\rm{element with highest mass number}}{\rm{.}}\\\end{array}

If the higher priority groups are on the opposite side of the double bonded carbon atoms, then it is designated as E.

If the higher priority groups are on the same side of the carbon atoms, then it is designated as Z.

It has 9 carbon atoms in the longest carbon chain. An ethyl group is present at the fifth position as substituent. It has two double bonds at the carbon atoms 3 and 5.

The structure of (3E,5Z)5methyl3,5nonadiene\left( {3E,5Z} \right){\rm{ - 5 - methyl - 3,5 - nonadiene}} is as follows:

CH3
CH2
7
Н.С
На
с.
Н,С 2
С,
4
С
н
CH
6
3
1
Н,с.
CH3
HC

Ans:

The structure of (3E,5Z)5methyl3,5nonadiene\left( {3E,5Z} \right){\rm{ - 5 - methyl - 3,5 - nonadiene}} is as follows:

CH3
CH2
7
Н.С
На
с.
Н,С 2
С,
4
С
н
CH
6
3
1
Н,с.
CH3
HC